OpenStudy (anonymous):

Square QRST has vertices q(2,8), r(-4,8), s(-4,2), and t(2,2),and triangle QTU has vertices q(2,8), t(2,2), and u(6,2). What is the area of the quadrilateral formed by the square and triangle

OpenStudy (anonymous):

Please help

OpenStudy (nurali):

One side of the square measures 6 units. Q(2,8) and R (-4, 8) The y-values are the same, so I added the absolute values of 2 and -4. |2|= 2 |-4| = 4 2 + 4= 6 Area of square QRST: 6 x 6 = 36 sq. units The height of the triangle is 6 and the base is 4 units. TRIANGLE: Area of Triangle QTU: 6 x 4 / 2= 12 sq. units 36 + 12 = 48 sq. units

OpenStudy (anonymous):

Thank you

OpenStudy (nurali):

My Pleasure.